1. x+4y=14
2x-y=1 solve using substitution method

Answer by rm29924(97)

take the first equation and get x by itself
x+4y = 14
to get x by itself subtract 4y from both sides
x+4y = 14
-4y -4y
x= 14-4y
now substitue this value for x in the second equation
2x-y=1
2(14-4y)-y=1
multiply everything in the parenthesis by 2
2(14) 2(-4y)
28-8y-y=1
28-9y=1
add 9y to both sides
28=1+9y
subtract 1 from both sides
27=9y
divide both sides by 9
3=y
now substitue the value for y in the first equation to get x
x+4y=14
x+4(3)=14
x+12=14
subtract 12 from both sides
x=2
answer: x=2 y=3
check work:
x+4y=14
2+4(3)=14
2+12=14
14+14
2x-y=1
2(2)-3=1
4-3=1
1=1